OpenStudy (anonymous):

Verify that the functions f and g , are inverses of each other by showing that f(g(x))=x and g(f(x))=x . Graph both functions on the same graph. f(x)-(x+2)^3 ; g(x)=^3 sqrt x-2

OpenStudy (anonymous):

f(x) = (x+2 )^3 and g(x) = ?

OpenStudy (anonymous):

g(x)= exponent 3 then sqrt x-2

OpenStudy (anonymous):

\[\sqrt[3]{x-2}?\]

OpenStudy (anonymous):

^ yes that.

OpenStudy (anonymous):